Hastings Corporation is interested in acquiring Vandell Corporation. Vandell has 1 million shares outstanding and a target capital structure consisting of 30% debt; its beta is 1.40 (given its target capital structure). Vandell has $10.71 million in debt that trades at par and pays an 7.5% interest rate. Vandell’s free cash flow (FCF0) is $1 million per year and is expected to grow at a constant rate of 5% a year. Both Vandell and Hastings pay a 35% combined federal and state tax rate. The risk-free rate of interest is 4% and the market risk premium is 8%. Hastings Corporation estimates that if it acquires Vandell Corporation, synergies will cause Vandell’s free cash flows to be $2.4 million, $3.2 million, $3.5 million, and $3.52 million at Years 1 through 4, respectively, after which the free cash flows will grow at a constant 5% rate. Hastings plans to assume Vandell’s $10.71 million in debt (which has an 7.5% interest rate) and raise additional debt financing at the time of the acquisition. Hastings estimates that interest payments will be $1.5 million each year for Years 1, 2, and 3. After Year 3, a target capital structure of 30% debt will be maintained. Interest at Year 4 will be $1.463 million, after which the interest and the tax shield will grow at 5%. Indicate the range of possible prices that Hastings could bid for each share of Vandell common stock in an acquisition. Round your answers to the nearest cent. Do not round intermediate calculations. The bid for each share should range between $ per share and $ per share.
Answer #1

Weighted Cost of Capital for Vandell corporation

W = W_d*C_d+W_e*C_e

cost of debt (Cd) = I (1-T) = 7.5 *(1-0.35) = 4.875%

Cost of Equity (Ce) = Risk free + Beta * Market Risk premium = 4 + 1.4*8 = 15.2%

W = 0.30*4.875% + 0.7*15.2% =12.103%

Value of Vandell with synergy by use of discounting cash flow approach .

Value of firm (V) = Cash flow Cash flow2 Cash flow3 Cash flows Cash flow (1+g)

V = 2.4 3.2 3.5 3.52 3.52(10.05) (10.12103) (1.12103) 1.12103)3 (1.12103)4 (0.12103 0.05 (1.12103)4

V = $91.579 millions

if 30% portion is debt , so value of equity = $ 64.105 million

Value per share = $ 64.105 = $64.11

Similarly the value of debt raised to acquire the company :

Value of debt = Cash flow Cash flow2 Cash flow3 Cash flows Cash flow (1+g)

value of debt = 1.5 1.5 1.463 1.463 * (1 0.05) 1.5 (10.12103) (1.12103) 1.12103)3 (1.12103)4 (0.12103 0.05 (1.12103)4

Value of debt =$ 38.678 millions

Net value of firm = 91.578 - 38.678 = $ 52.899 million

Value of share = $ 52.899 = $52.90

Bid of share should range from $ 52.90 to  $ 64.11